SakunasWifeu:

What is 17 1/1 percent of $1800

SakunasWifeu:

1/2*

boredfr:

17.5% of $1800 17.5 percent *1800 = (17.5:100)*1800 = (17.5*1800):100 = 31500:100 = 315

SakunasWifeu:

Thank uuu

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SakunasWifeu Thank uuu \(\color{#0cbb34}{\text{End of Quote}}\) Yw

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SakunasWifeu Im dyibg cuz of math rn :'( \(\color{#0cbb34}{\text{End of Quote}}\) Well if you have any other questions like this just tag me

SakunasWifeu:

Ahh okk thank uuu

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SakunasWifeu Ahh okk thank uuu \(\color{#0cbb34}{\text{End of Quote}}\) hmmm np

SakunasWifeu:

@boredfr can you tell me 3.5 percent of 7035

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SakunasWifeu @boredfr can you tell me 3.5 percent of 7035 \(\color{#0cbb34}{\text{End of Quote}}\) 3.5 percent *201000 = ( 3.5:100)*201000 = ( 3.5*201000):100 = 703500:100 = 7035

boredfr:

Close this post and make a new one for the next on and just tag me in that one

SakunasWifeu:

Kk